# A dialog between cell adhesion and topology at the core of morphogenesis

### v1

@@ Acknowledgements.
We thank Dimitri Fábreges, Takashi Hiiragi and Nicoletta I. Petridou for comments and generously sharing imaging pictures of in vivo embyonic tissues. This work was supported by the Weave project ”Tissue material phase transitions and their role in embryo pattern formation” from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, 518354236, PE 3800/1-1) and Österreichischer Wissenschaftsfonds (FWF, Austrian Science Fund, I6533) (E. F and B. C-M). B.C-M. and A. A-T. acknowledge the support of the field of excellence “Complexity of life in basic research and innovation” of the University of Graz.

### v2

@@ Acknowledgements.
We thank Dimitri Fábreges, Takashi Hiiragi and Nicoletta I. Petridou for comments and generously sharing imaging pictures of in vivo embyonic tissues. This work was supported by the Weave project ”Tissue material phase transitions and their role in embryo pattern formation” from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, 518354236, PE 3800/1-1) and Österreichischer Wissenschaftsfonds (FWF, Austrian Science Fund, I6533) (E. F and B. C-M). B.C-M. and A. A-T. acknowledge the support of the field of excellence ”Complexity of life in basic research and innovation” of the University of Graz.